OpenStudy (anonymous):

Find the equation of the tangent plane to the surface z=4x2+8y2+9xy at the point (3,−3,27).

OpenStudy (anonymous):

Not really sure what i'm missing or doing wrong. fx=8x+9y, fy=16y+9x. fx(3,-3)=8*3+9*-3=-3 and fy(3,-3)=16*.3+9*3=-21. so should have z-27=8(x+3)+9y+16(y+21)+9x